OpenStudy (anonymous):

Help Please! Will Give Fan, Medal and Testomonies. Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B has a volume of 3456 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

@tinybookworm

OpenStudy (anonymous):

i know how to do it and get the answer i just dont know how to put it into a percentage

OpenStudy (anonymous):

other openstudy forums say to jest multiply the number by 100 but thats not what i learned in the lesson. I just cant remember exactly what i learned.

OpenStudy (anonymous):

@geerky42 can you help?

OpenStudy (anonymous):

@Mathhelp346

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@jane11509

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@magepker728

OpenStudy (studygurl14):

First, find the volume of Pyramid A. Do you know the formula for volume of a pyramid?

OpenStudy (anonymous):

yeah. I got that. its v=1/3bh

OpenStudy (anonymous):

height is 8 base length is 12. v=1/3(12)(8) v=1/3(96) v=32

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

How about 3456 divide by 32 time 100?

OpenStudy (anonymous):

then i would get 10800%

OpenStudy (anonymous):

Yes